Project

General

Profile

10/24/2013 Minutes

Attendance

Joao, Susan, Nick, Mat, Gavin, Athans, Dom, Kanika, Evan

Operations

Susan reports:
  • APD work going on (running with empty runs)
  • 8 di-blocks running cold overnight (some problem)

Production

  • We have a branch tag B13-10-23.
    1. Gavin and Dom will work it out
  • NuMI caf Numi not showing up.
  • From run 11403 1/10 events of cosmic triggers being processed.
    1. Missing files resolved by Robert Illingsworth. Bugs in FTS. Files now appear at FNAL
  • 'find' command processes most recent files first now

Michael showed some quick slides showing that slicer chops tracks on electronics sampling boundaries timing. Fully report tomorrow. Will produce multiple parameter sets for event scanning.

Kanika reports that the Nearline looked good the last days

Nick reports that the data appears unmangled in NDOS. How much time needed to run on FD? Likely report at tomorrow's DQ meeting.

Outstanding Issues

  • Missing runs now rescued from Ash River and on disk at FNAL
  • Information missing from Logbook for about a third of the runs
  • Mat will look to see which files are at Ash River and on Bluearc. Check the consistency